OpenStudy (samigupta8):

the value of log x+ log(1+1/x)+log(1+1/1+x)+log(1+1/2+x)+.............+log(1+1/(n-1+x))

OpenStudy (amistre64):

i assume that not 1+1 over the rest

OpenStudy (samigupta8):

it's only under 1 not whole

OpenStudy (amistre64):

the addition of logs refers to multiplication of arguments \[log(a)+lob(b)=log(ab)\]

OpenStudy (amistre64):

the division of logs referes to subtraction ... so i believe there is a telesoping part to thos

OpenStudy (amistre64):

log x +log (1+x)/(x) +log (x+2)/(1+x) +log(3+x)/(2+x) +............. +log(n+x)/(n-1+x) log (x) +log (1+x) -log(x) +log (x+2) -log(1+x) +log(3+x) -log(2+x) +............. +log(n+x)/(n-1+x) log(n+x) this sums out to

OpenStudy (raden):

or you could simplify of the terms : (1+1/x) = (x+1)/x 1+1/(1+x) = (x+2)/(x+1) 1+1/(2+x) = (x+3)/(x+2) 1+1/(3+x) = (x+4)/(x+3) ... ... 1+1/(n-1+x) = (x+n)/(x+n-1) then use the property of log, it becomes